Calculate the mole fraction of the solution in a solution containing 85g ethanol in 500g of


water

1
Expert's answer
2021-11-24T05:42:01-0500

The question requirements is to get the mole fraction of the solution


"\Chi_{solution}=1"


But adjusting the condition of the question to read;


" Calculate the mole fraction of each component of a solution containing "85g" ethanol in "500g" of water"




Solution



"\Chi_A=\frac{n_A}{n_A+n_B}"


Where "\Chi_A" is mole fraction of A

"n_A=" Moles of A

"n_B=" Moles of B



"n_{water}=\frac{500}{18}=27.7778"


Molar mass of ethanol "=46g\/mol"


"n_{ethanol}=\frac{85}{46}=1.8478"


Total moles "=27.7778+1.8478=29.6256"


"\Chi_{water}=\frac{27.7778}{29.6256}=0.9376"


"\Chi_{ethanol}=\frac{1.8478}{29.6256}=0.0624"